water vapor changed from 100c to 25c in a 12 oz bottle. what is final volume?
Answer #1

Answer:

Initial volume = 12 oz = V1

Let final volume = V = V2

Initial temperature = T1 = 100oC = 373 K

Final temperature = 25oC = 298 K

For a gaseus system, V/T is a constant quantity,

in other words

V1/T1 = V2/T2

So V2 = (V1/T1) x T2 = (12 Oz / 373 K) x 298 K = 9.6 Oz

So final volume of water vapor is 9.6 Oz.
Answer #1

Answer:

Initial volume = 12 oz = V1

Let final volume = V = V2

Initial temperature = T1 = 100oC = 373 K

Final temperature = 25oC = 298 K

For a gaseus system, V/T is a constant quantity,

in other words

V1/T1 = V2/T2

So V2 = (V1/T1) x T2 = (12 Oz / 373 K) x 298 K = 9.6 Oz

So final volume of water vapor is 9.6 Oz.